OpenStudy (skullcandy229):

If A = -3n + 2, and B = 5n - 7, what is the value of A + B, in simplest form? I will give you a medal

OpenStudy (solomonzelman):

EXAMPLE: You are given that: \(\large\color{green}{ \displaystyle A = 4n - 1 }\) \(\large\color{blue}{ \displaystyle B = -6n+2 }\) \(\large\color{black}{ \displaystyle \color{green}{A}+\color{blue}{B} = }\) \(\large\color{black}{ \displaystyle \color{green}{(4n - 1)}+\color{blue}{( -6n+2)} = }\) \(\large\color{black}{ \displaystyle 4n - 1-6n+2= }\) \(\large\color{black}{ \displaystyle 4n -6n+2- 1= }\) \(\large\color{black}{ \displaystyle -2n+1. }\)